# Best Practices for Biorisk Evaluations on Open-Weight Bio-Foundation Models

### v3

@@ B.1 Hardware Configuration
We use Amazon p5.48xlarge [^5] https://aws.amazon.com/ec2/instance-types/p5/ as our experiment platform, which consists of NVIDIA H100-80GB GPUs and AMD EPYC 7R13 Processor. All the experiments (fine-tune and inference) are done with 4 NVIDIA H100-80GB GPUs under NVIDIA BioNeMo Framework [25].
@@ Appendix C Evo 2 representation Analysis
In this section, we discuss the architecture and Evo 2 and explain why we observe extremely large representations in the later layers, but can still get reasonable output. Evo 2 uses StripedHyena2 [28] architecture, which builds blocks from input-dependent long convolutions. A single Hyena operator creates three streams $q,k,v$ by convolving linear projects of the residual stream $x$ with Toeplitz filters $T,H,K$, then mixes them with inner convolution $G$ and a final projection $M$ to get the output $y$:

### v4

@@ B.1 Hardware Configuration
We use Amazon p5.48xlarge [^5] https://aws.amazon.com/ec2/instance-types/p5/ as our experiment platform, which consists of NVIDIA H100-80GB GPUs and AMD EPYC 7R13 Processor. All the experiments (fine-tune and inference) are done with 4 NVIDIA H100-80GB GPUs under NVIDIA BioNeMo Framework [john2024bionemo].
@@ Appendix C Evo 2 representation Analysis
In this section, we discuss the architecture and Evo 2 and explain why we observe extremely large representations in the later layers, but can still get reasonable output. Evo 2 uses StripedHyena2 [ku2025systems] architecture, which builds blocks from input-dependent long convolutions. A single Hyena operator creates three streams $q,k,v$ by convolving linear projects of the residual stream $x$ with Toeplitz filters $T,H,K$, then mixes them with inner convolution $G$ and a final projection $M$ to get the output $y$: